# SkyDiffusion: Street-to-Satellite Image Synthesis with Diffusion Models and BEV Paradigm

### v2

@@ Abstract
Street-to-satellite image synthesis focuses on generating realistic satellite images from corresponding ground street-view images while maintaining a consistent content layout, similar to looking down from the sky. The significant differences in perspectives create a substantial domain gap between the views, making this cross-view generation task particularly challenging. In this paper, we introduce SkyDiffusion, a novel cross-view generation method for synthesizing satellite images from street-view images, leveraging diffusion models and Bird’s Eye View (BEV) paradigm. First, we design a Curved-BEV method to transform street-view images to the satellite view, reformulating the challenging cross-domain image synthesis task into a conditional generation problem. Curved-BEV also includes a "Multi-to-One" mapping strategy for leveraging multiple street-view images within the same satellite coverage area, effectively solving the occlusion issues in dense urban scenes. Next, we design a BEV-controlled diffusion model to generate satellite images consistent with the street-view content, which also incorporates a light manipulation module to make the lighting conditions of the synthesized satellite images more flexible. Experimental results demonstrate that SkyDiffusion outperforms state-of-the-art methods on both suburban (CVUSA & CVACT) and urban (VIGOR-Chicago) cross-view datasets, with an average SSIM increase of 13.96% and a FID reduction of 20.54%, achieving realistic and content-consistent satellite image generation. The code and models of this work will be released at https://opendatalab.github.io/skydiffusion/
Introduction
Cross-view synthesis involves transformations between disparate viewpoints, such as from street-view images to satellite images and vice versa. This process is crucial for applications in urban planning (Li et al. 2023b), 3D reconstruction (Wang et al. 2021), virtual reality (Emmaneel et al. 2023), etc. The significant changes in street view and satellite perspectives result in drastically different appearances of objects, leading to a substantial disparity between different image domains. As shown in Fig 1 (a), in cross-view scenarios, building facades observed only in street view and rooftops observed only from satellite view are considered uncorrelated and drastically different features (Regmi and Borji 2018; Zhu et al. 2023). However, objects such as roads and lawns, which are below the viewpoint, can be observed in different views even if they undergo some deformation, making them cross-view shared correlated information. The key to the challenging task of cross-view synthesis lies in bridging domain gaps, capturing cross-view relevant information, and generating realistic views with consistent content layouts.
Figure 1: Illustration of the cross-view image synthesis task. (a) In the cross-view scenario, there are features unique to the street-view image, features exclusive to the satellite image, and shared correlated information observable from both views. (b) Comparison of our street-to-satellite image synthesis method with existing cross-view synthesis methods (e.g., SelectionGAN (Tang et al. 2019)) or image translation methods based on diffusion models (e.g., Instr-p2p (Brooks, Holynski, and Efros 2023)).
In the field of cross-view synthesis, converting satellite images to street-view images has received extensive attention (Shi et al. 2022; Lu et al. 2020; Qian et al. 2023). Researchers have leveraged polar transformations (Toker et al. 2021) or geo-transformation (Lu et al. 2020) to bridge the gap from the satellite to the street view domain, and then employ Generative Adversarial Networks (GANs) to generate realistic street views. This paper focuses on the less explored yet equally challenging task of synthesizing satellite images from street-view images. Some researchers have used semantic segmentation priors to aid in synthesizing satellite images, achieving notable results (Tang et al. 2019). However, the semantic maps are often unavailable in real scenarios. Existing street-to-satellite image generation also lacks transformations such as polar or geo-transformation to fit the gap from the street view to the satellite domain. Moreover, as shown in Fig 1(b), existing methods on street-to-satellite image synthesis, which mostly rely on GANs, often result in lower-quality satellite images with unrealistic textures.
Recently, diffusion models have gained attention for their exceptional image synthesis (Ho, Jain, and Abbeel 2020; Zhang, Rao, and Agrawala 2023). Many studies have explored text-guided fine-tuning of diffusion models to top-down perspective from other views (Kothandaraman et al. 2023a; Ruiz et al. 2023). Additionally, novel view synthesis methods have been developed by training with multi-view images and camera poses (Liu et al. 2023; Shi et al. 2023a). Although current novel view synthesis techniques demonstrate satisfactory performance at the object level, their performance in cross-view generation for complex scenes with significant viewpoint differences remains suboptimal. A possible solution is to treat the this task as an image translation task (Li et al. 2023a; Parmar et al. 2024). However, as shown in Fig 1(b), due to the domain gap of different views, the resulting images are often realistic but not actual, with significant discrepancies between the synthesized satellite images and the actual corresponding street view content. In response to these challenges, this paper introduces SkyDiffusion, a street-to-satellite image synthesis method based on diffusion models and the BEV paradigm, similar to viewing from the sky. Unlike previous cross-view synthesis methods, our approach is based on a diffusion framework, enabling the generation of more realistic satellite images. We design a Curved-BEV method that transforms street-view images into satellite views, achieving domain alignment between different perspectives and facilitating content-consistent satellite image synthesis. To address the limited perception range of BEV in densely occluded urban scenarios, the Curved-BEV method also includes a Multi-to-One mapping strategy. Additionally, our BEV-controlled diffusion model incorporates a light manipulation module, making the lighting conditions of the synthesized satellite images more adaptable. Our main contributions are summarized as follows:
@@ Introduction
  • We introduce SkyDiffusion, a novel street-to-satellite image synthesis method leveraging diffusion models and BEV paradigm to generate realistic and content-consistent satellite images.
@@
  • SkyDiffusion outperforms state-of-the-art methods on suburban (CVUSA, CVACT) and urban (VIGOR-Chicago) cross-view datasets, with a 13.96% SSIM increase and 20.54% FID reduction.
Related Work
@@
Cross-view image synthesis involves generating realistic scenes images with significant viewpoint changes, primarily divided into satellite-to-street (Shi et al. 2022; Lu et al. 2020; Qian et al. 2023) and street-to-satellite generation (Regmi and Borji 2018, 2019). In this field, synthesizing satellite images from street views is a challenging yet important task. Regmi and Borji (2018) attempted to synthesize aerial views from a single street image, but the significant domain differences led to inconsistencies. Tang et al. (2019) used street-view images and aerial semantic maps for image-to-image translation to synthesize target satellite images, but semantic maps are often unavailable in real scenarios. Additionally, the performance of GANs has also limited the quality of street-to-satellite image synthesis in previous studies.
@@
Figure 2: Overview of the proposed SkyDiffusion framework, including the curved BEV transformation and BEV-controlled diffusion model with light manipulation.
@@
BEV transformation is a perspective domain conversion technique similar to polar transformation (Shi et al. 2019; Toker et al. 2021) and geo-transformation (Lu et al. 2020), widely used in retrieval (Ye et al. 2024b), navigation (Li et al. 2022) and localization tasks (Song et al. 2023; Wang et al. 2023; Shi et al. 2023b). Traditional BEV transformation techniques, which rely on camera parameters (Sarlin et al. 2023) or depth estimation methods (Teng et al. 2024; Ye et al. 2024a), face challenges such as the difficulty in obtaining camera parameters and the high computational cost of depth estimation. Recently, some methods have been proposed to map street-view images to the BEV perspective using geometric relations at the image (Wang et al. 2023) or feature level (Shi et al. 2023b). However, such methods are based on the ground plane assumption, where all mapped points are located below the camera, limiting the mapping of upper street scene features and causing information loss. Furthermore, to our knowledge, no studies have yet explored the use of BEV methods for street-to-satellite image synthesis task.
@@
Novel view synthesis based on diffusion models. Diffusion models have shown significant potential in recent studies on novel view synthesis (Kawar et al. 2023; Gao et al. 2024; Tseng et al. 2023). DreamBooth (Ruiz et al. 2023) and AerialDiffusion (Kothandaraman et al. 2023a) utilize text guidance to achieve top-down image generation by fine-tuning diffusion models. However, it is challenging to fully encapsulate the complexity of real-world scenes using textual information alone. Zero-1-to-3 (Liu et al. 2023) and Zero 123++ (Shi et al. 2023a) use multi-view images and camera poses to guide diffusion models in learning object structures and textures, enabling new viewpoint synthesis. However, these methods are limited to object-level tasks with strong viewpoint correlations and struggle to handle the significant viewpoint changes in cross-view generation. Besides, diffusion-based image-to-image translation methods can also generate realistic satellite images (Li et al. 2023a; Brooks, Holynski, and Efros 2023; Zhang, Rao, and Agrawala 2023). However, the substantial domain gap between street-view images and satellite images presents a significant challenge in generating content-consistent satellite images.
Methods
@@
As illustrated in Fig. 2, this paper introduces SkyDiffusion, a novel method for synthesizing satellite images from corresponding street-view images. It initially applies a curved BEV transformation to the input street-view images, converting the perspective to a top-down view using either one-to-one or multi-to-one mappings. Subsequently, the BEV-Controlled diffusion model with light manipulation is employed for the controlled synthesis of satellite images. Following this, we detail our network training process.
Curved BEV Transformation
@@
One-to-One BEV. Existing BEV methods based on ground plane assumptions ($P(x,y,z=0)$) and geometric relationships achieve the mapping from street view to the BEV plane(Shi et al. 2023b; Wang et al. 2023). However, this assumption fails to map the content above the street-view image and results in noticeable distortion as points move away from the center area. Thus, we propose an enhanced curved BEV transformation to better capture cross-view information and address this issue.
Figure 3: Schematic of the curved BEV transformation. The upper part (a) details the transformation process, highlighting the two mapping points at the top and bottom of the street view. The lower parts (b) and (c) present the results of One-to-One and Multi-to-One BEV transformations, respectively.
@@
As shown in Fig. 3 (a), we use $P(i,j)$ to denote an index on the BEV space, and $P(x,y,z)$ to illustrate a point in a three-dimensional coordinate system with the camera on the z-axis. We assume that the BEV plane is an upward curved surface, with the height (z-axis) increasing rapidly as the 3D points move away from the center (the central part of Fig 3(a)). The following equation formulates the relationship between $P(x,y,z)$ and $P(i,j)$, where $d_{norm}$ represents the normalized distance from the center, $d_{max}$ is the maximum distance from the center, and $\lambda$ is a scaling factor.
@@
Furthermore, we can convert $P(x,y,z)$ to spherical coordinates $P(\theta,\varphi)$. Based on the equirectangular projection properties of street view images, $P(\theta,\varphi)$ corresponds to the panorama image index $P(u,v)$. This allows us to establish the mapping relationship between $P(x,y,z)$ and $P(u,v)$ from the following equation, where $H$ represents the camera height, and $w$, $h$ represent the width and height of the street image. More details are in the supplementary material.
@@
Following Eq. 1 and Eq. 2, we achieve the projection from the street-view image to the bird’s eye view without requiring camera calibration parameters or depth estimation methods (Teng et al. 2024; Ye et al. 2024a) (Fig. 3 (b)). This effectively maps upper street view features like distant roads and buildings. Since the mapping relationship for each transformation is fixed, the computational overhead is minimal.
Multi-to-One BEV. Due to the limitations imposed by dense buildings on street view observations, the BEV sensing range is very limited in urban datasets like VIGOR. To address this, we propose a "Multi-to-one" BEV mapping method that unifies multiple street view BEV mappings results $\textit{BEV}_{\mathrm{cam}_{k}}$ into a single satellite coordinate system based on the relationships of camera position (Fig. 3 (c)). We use a simple semantic segmentation network to identify and discard the regions with severe BEV distortion, typically those obstructed by buildings. We select the street view mapping result closest to the capture point for overlapping areas following Eq. 3, where $\textit{BEV}\left(x,y\right)$ represents the unified BEV space corresponding to the satellite image, $x,y$ corresponds to the index position of the BEV plane, $\Delta x_{k}$ and $\Delta y_{k}$ represent the offset of the street-view image from the satellite center, and $x_{\mathrm{cam}_{i}}$ and $y_{\mathrm{cam}_{i}}$ represent the coordinates of camera position in the satellite coordinate system.
@@
BEV-Controlled Synthesis with Light Manipulation
Curved BEV controlled synthesis. Image diffusion models progressively denoise images and generate samples from the training domain. The denoising process can occur in pixel space or in a latent space encoded from training data. Inspired by ControlNet (Zhang, Rao, and Agrawala 2023), we enhance the control over the generated aerial views by replicating the SD module, enabling the transformed BEV images to effectively guide the synthesis process. Specifically, we first use a small encoder to capture the cross-view associative information of the input BEV image $I_{bev}$ such as road directions and building positions, to achieve BEV image feature embedding $c_{bev}$. Notably, considering the potential feature distortions in BEV images, we employ a spatial attention mechanism to enhance feature capture while suppressing distorted features.
@@
Light Manipulation. Given the differences in perspectives and timing between satellite and street images, there is no direct correlation in lighting conditions. To address this, we designed a light manipulation module to make the lighting conditions of the synthesized satellite images more adaptable. We employ the DNCM for modeling deterministic color mapping (Ke et al. 2023). Initially, we use nDNCM to obtain the "image content information" $I_{c}$ from the preliminary synthesis results $I_{0}$ of the satellite images. Then, we input the lighting and color style $r_{s}$ obtained from the reference aerial view into sDNCM for lighting transfer.
@@
Figure 4: Qualitative Comparison of Synthesis Results. The first and second rows present the results on the CVUSA dataset, while the third and fourth rows display the results on the CVACT dataset. The fifth and sixth rows showcase the visualized outcomes on the VIGOR-Chicago dataset.
@@
Network Training
Loss Function. We train the BEV-controlled synthesis of satellite images and the light manipulation module in two separate stages. Given a set of cross-view image data, the image diffusion algorithm gradually adds noise to the satellite image, generating a noisy satellite image $Z_{t}$. It then learns a network $\epsilon_{\theta}$ to predict the noise added to the noisy satellite image $Z_{t}$ based on the given BEV image as the conditional input. Since text cannot fully describe the street view scene, our cross-view image synthesis is essentially an image-conditioned synthesis task, with text not being our optimization target. This helps the model focus on generating satellite images based on the BEV conditions.

### v3

# SkyDiffusion: Ground-to-Aerial Image Synthesis with Diffusion Models and BEV Paradigm

@@ Abstract
Ground-to-aerial image synthesis focuses on generating realistic aerial images from corresponding ground street view images while maintaining consistent content layout, simulating a top-down view. The significant viewpoint difference leads to domain gaps between views, and dense urban scenes limit the visible range of street views, making this cross-view generation task particularly challenging. In this paper, we introduce SkyDiffusion, a novel cross-view generation method for synthesizing aerial images from street view images, utilizing a diffusion model and the Bird’s-Eye View (BEV) paradigm. The Curved-BEV method in SkyDiffusion converts street-view images into a BEV perspective, effectively bridging the domain gap, and employs a "multi-to-one" mapping strategy to address occlusion issues in dense urban scenes. Next, SkyDiffusion designed a BEV-guided diffusion model to generate content-consistent and realistic aerial images. Additionally, we introduce a novel dataset, Ground2Aerial-3, designed for diverse ground-to-aerial image synthesis applications, including disaster scene aerial synthesis, historical high-resolution satellite image synthesis, and low-altitude UAV image synthesis tasks. Experimental results demonstrate that SkyDiffusion outperforms state-of-the-art methods on cross-view datasets across natural (CVUSA), suburban (CVACT), urban (VIGOR-Chicago), and various application scenarios (G2A-3), achieving realistic and content-consistent aerial image generation. More result and dataset information can be found at https://opendatalab.github.io/skydiffusion/.
1 Introduction
In fields such as land cover classification [18, 36], urban planning [47, 45], and disaster response [11, 44, 38], the analysis of aerial imagery plays a crucial role. The cross-view Ground-to-Aerial Image Synthesis task leverages ground street-view images to synthesize corresponding aerial images at the same location. Street-view images can be captured and uploaded by users on platforms such as Google Street View or Mapillary, offering high flexibility in data collection and frequent temporal coverage [1]. Therefore, Ground-to-Aerial Image Synthesis can support rapid response during natural disasters [19] or fill historical gaps in high-resolution satellite imagery records [3], among other applications.
Figure 1: Illustration of the cross-view image synthesis task. (a) Challenges of cross-view domain gaps; (b) Challenges of occlusion in dense scenes; (c) Comparing our ground-to-aerial image synthesis method with existing cross-view synthesis methods.
Despite the potential of cross-view synthesis from ground-level to aerial imagery, significant challenges remain. The drastic viewpoint differences between street and satellite views lead to distinct object appearances, resulting in substantial domain gaps across image types. As shown in Figure 1 (a), in cross-view scenarios, street-view images primarily capture ground and building facade details, while satellite views present more rooftop and macroscopic summary information. Even with state-of-the-art diffusion methods, the domain gap between views often results in generated images that are realistic yet not true to reality; the synthesized satellite imagery may differ significantly from the actual content of corresponding street-views, as shown in Figure 1(c).
Another challenge is object occlusion in dense urban settings, where tall buildings or trees limit the visibility range in street-views, which is substantially smaller than the coverage of satellite images, as shown in Figure 1(b). Aerial images synthesized from a single street-view image struggle to maintain consistency with real satellite images in unseen areas. A effective approach is to leverage multiple neighboring street-view images for satellite view synthesis. However, this poses a challenge for the model in effectively utilizing geographic location information from street-views and integrating features from multiple street-view images.
To address these challenges, this paper introduces SkyDiffusion, a ground-to-aerial image synthesis method based on diffusion models and the BEV paradigm, emulating a sky-view perspective. We first propose a Curved-BEV method, which transforms street-view images into Bird’s-Eye view to achieve domain alignment across viewpoints. Additionally, the Curved-BEV method introduces a multi-to-one mapping strategy, effectively integrating street-view images from different capture locations into a unified BEV space, thus overcoming the limited observation range in urban street-view images. Subsequently, we designed a BEV-guided diffusion model to generate realistic satellite images consistent with street-view content. Therefore, SkyDiffusion effectively overcomes the challenges of cross-view domain gaps and the limited observation range in densely occluded urban scenes, generating more realistic satellite images.
Moreover, existing cross-view datasets, such as CVUSA [40] and VIGOR[46], were originally designed for cross-view retrieval tasks and thus offer limited diversity for cross-view synthesis tasks. We introduce an additional dataset, Ground2Aerial-3 (G2A-3), which includes three novel ground-to-aerial scene synthesis tasks: aerial image synthesis for disaster scenes based on Florida hurricane data, high-resolution historical aerial image generation (2007-2014), and low-altitude UAV image generation based on the MatrixCity [14] virtual urban environment. Ground2Aerial-3 provides practical value and significance for exploring cross-view synthesis tasks across more diverse scenarios.
Our main contributions are as follows:
@@ 1 Introduction
  • We introduce SkyDiffusion, a novel ground-to-aerial synthesis method leveraging diffusion models and BEV paradigm to generate realistic, consistent aerial images.
@@
  • We introduce Ground2Aerial-3, a new ground-to-aerial image synthesis dataset, featuring disaster scene aerial image, historical high-resolution satellite image, and low-altitude UAV image synthesis tasks.
@@
  • SkyDiffusion outperforms state-of-the-art methods on cross-view datasets across natural (CVUSA), suburban (CVACT), urban (VIGOR-Chicago), and various application scenarios (G2A-3), with an SSIM increase of 8.67% and a FID reduction of 21.50%.
2 Related Work
@@
Cross-view image synthesis involves generating realistic scenes images with significant viewpoint changes, primarily divided into aerial-to-ground [29, 17, 22] and ground-to-aerial generation [23, 24]. In this field, synthesizing satellite images from street-views is a challenging yet important task. Regmi and Borji [23] attempted to synthesize aerial views from a single street image, but the significant domain differences led to inconsistencies. Tang et al. [33] used street-view images and aerial semantic maps for image-to-image translation to synthesize target satellite images, but semantic maps are often unavailable in real scenarios. Additionally, the performance of GANs has also limited the quality of ground-to-aerial image synthesis in previous studies. Concurrent work [1] also synthesizes satellite images using text guidance and staged BEV semantic segmentation results, but these methods require additional text data and an extra semantic segmentation phase on street-view images.
@@
Figure 2: Overview of the proposed SkyDiffusion framework. It include the curved BEV transformation and BEV-controlled diffusion model with light manipulation. The lower parts present the results of One-to-One and Multi-to-One BEV transformations, respectively.
@@
BEV transformation is a perspective domain conversion technique similar to polar transformation [28, 35] and geo-transformation [17], widely used in retrieval [42], navigation [15] and localization tasks [32, 39, 30]. Traditional BEV transformation techniques, which rely on camera parameters [27] or depth estimation methods [34, 41], face challenges such as the difficulty in obtaining camera parameters and the high computational cost of depth estimation. Recently, some methods have been proposed to map street-view images to the BEV perspective using geometric relations at the image [39] or feature level [30]. However, such methods are based on the ground plane assumption, where all mapped points are located below the camera, limiting the mapping of upper street scene features and causing information loss. Moreover, most existing panoramic BEV transformation methods primarily focus on one-to-one mappings, with relatively little research on multi-to-one BEV joint mappings.
@@
Novel view synthesis based on diffusion models. Diffusion models have shown significant potential in recent studies on novel view synthesis [7, 4, 37]. DreamBooth [26] and AerialDiffusion [9] utilize text guidance to achieve top-down image generation by fine-tuning diffusion models. However, it is challenging to fully encapsulate the complexity of real-world scenes using textual information alone. Besides, diffusion-based image-to-image translation methods can also generate realistic satellite images [12, 2, 43]. However, the substantial domain gap between street-view images and satellite images presents a significant challenge in generating content-consistent satellite images.
3 Methods
@@
As illustrated in Fig. 2, this paper introduces SkyDiffusion, a novel method for synthesizing satellite images from corresponding street-view images. It initially applies a curved BEV transformation to the input street-view images, converting the perspective to a top-down view using either one-to-one or multi-to-one mappings (Section 3.1). Subsequently, a BEV-guided diffusion model controls the synthesis of satellite images, with an optional light manipulation module included(Section 3.2). Following this, we detail our network training process (Section 3.3).
3.1 Curved BEV Transformation
One-to-One BEV.
@@ 3.1 Curved BEV Transformation
Existing BEV methods based on ground plane assumptions ($P(x,y,z=0)$) and geometric relationships achieve the mapping from street-view to the BEV plane[30, 39]. However, this assumption fails to map the content above the street-view image and results in noticeable distortion as points move away from the center area. Thus, we propose an enhanced curved BEV transformation to better capture cross-view information and address this issue.
@@
As shown in Fig. 3, we use $P(i,j)$ to denote an index on the BEV space, and $P(x,y,z)$ to illustrate a point in a three-dimensional coordinate system with the camera on the z-axis. We assume that the BEV plane is an upward curved surface, with the height (z-axis) increasing rapidly as the 3D points move away from the center (the central part of Fig 3(a)). The following equation formulates the relationship between $P(x,y,z)$ and $P(i,j)$, where $d_{norm}$ represents the normalized distance from the center, $d_{max}$ is the maximum distance from the center, and $\lambda$ is a scaling factor.
@@
Furthermore, we can convert $P(x,y,z)$ to spherical coordinates $P(\theta,\varphi)$. Based on the equirectangular projection properties of street-view images, $P(\theta,\varphi)$ corresponds to the panorama image index $P(u,v)$. This allows us to establish the mapping relationship between $P(x,y,z)$ and $P(u,v)$ from the following equation, where $H$ represents the camera height, and $w$, $h$ represent the width and height of the street image. More details are in the supplementary material.
@@
Following Eq. 1 and Eq. 2, we achieve the projection from the street-view image to the bird’s eye view without requiring camera calibration parameters or depth estimation methods [34, 41]. This effectively maps upper street-view features like distant roads and buildings. Since the mapping relationship for each transformation is fixed, the computational overhead is minimal.
Figure 3: Schematic of the curved BEV transformation. It illustrates the mapping of two points on the BEV plane to the top and bottom of the street-view image during transformation.
Multi-to-One BEV.
Due to the limitations imposed by dense buildings on street-view observations, the BEV sensing range is very limited in urban datasets like VIGOR. To address this, we propose a "Multi-to-one" BEV mapping method that unifies multiple street-view BEV mappings results $\textit{BEV}_{\mathrm{cam}_{k}}$ into a single satellite coordinate system based on the relationships of camera position. We select the street-view mapping result closest to the capture point for overlapping areas following Eq. 3, where $\textit{BEV}\left(x,y\right)$ represents the unified BEV space corresponding to the satellite image, $x,y$ corresponds to the index position of the BEV plane, $\Delta x_{k}$ and $\Delta y_{k}$ represent the offset of the street-view image from the satellite center, and $x_{\mathrm{cam}_{i}}$ and $y_{\mathrm{cam}_{i}}$ represent the coordinates of camera position in the satellite coordinate system.
@@
3.2 BEV-Controlled Diffusion Model
Curved BEV controlled synthesis. Image diffusion models progressively denoise images and generate samples from the training domain. The denoising process can occur in pixel space or in a latent space encoded from training data. Inspired by ControlNet [43], we enhance the control over the generated aerial views by replicating the SD module, enabling the transformed BEV images to effectively guide the synthesis process. Specifically, we first use a small encoder to capture the cross-view associative information of the input BEV image $I_{bev}$ such as road directions and building positions, to achieve BEV image feature embedding $c_{bev}$. Notably, considering the potential feature distortions in BEV images, we employ a spatial attention mechanism to enhance feature capture while suppressing distorted features.
@@ 3.2 BEV-Controlled Diffusion Model
Optional Light Manipulation. Given the differences in perspectives and timing between satellite and street images, there is no direct correlation in lighting conditions. To address this, we designed a light manipulation module to make the lighting conditions of the synthesized satellite images more adaptable. We employ the DNCM for modeling deterministic color mapping [8]. Initially, we use nDNCM to obtain the "image content information" $I_{c}$ from the preliminary synthesis results $I_{0}$ of the satellite images. Then, we input the lighting and color style $r_{s}$ obtained from the reference aerial view into sDNCM for lighting transfer.
@@
3.3 Network Training
Loss Function. We train the BEV-controlled synthesis of satellite images and the light manipulation module in two separate stages. Given a set of cross-view image data, the image diffusion algorithm gradually adds noise to the satellite image, generating a noisy satellite image $Z_{t}$. It then learns a network $\epsilon_{\theta}$ to predict the noise added to the noisy satellite image $Z_{t}$ based on the given BEV image as the conditional input. Since text cannot fully describe the street-view scene, our cross-view image synthesis is essentially an image-conditioned synthesis task, with text not being our optimization target. This helps the model focus on generating satellite images based on the BEV conditions.